IN WHAT PROPORTION SHOULD WATER BE MIXED IN A LIQUID OF RS 12 PER LITRE SO THAT WHEN ITS SOLD AT RS 13.75 PER LITRE , THERE IS PROFIT OF 25% PROFIT
1
Expert's answer
2012-12-06T09:16:29-0500
If the profit is 25%, getting:
13.75 = P(1 + 0.25), where P is the price of solution.
P = 13.75 / 1.25 = 11
Assuming that a liquid of RS 12 per litre is dissolved with a water of RS 0 per
litre. Then each litre of the solution costs:
(1 - X) * 12 + X * 0 = 11
So, 1 - X = 11 / 12
X / (1 - X) = (1 / 12) / (11 / 12) = 1 / 11.
Answer: 1 : 11.